# Dipolar ordering transitions in many-body quantum optics: Analytical diagrammatic approach to equilibrium quantum spins

### v2

@@ V.1 Long-range square lattice Heisenberg model
with $J_{1}<0$. This model has been recently studied with QMC simulations [42] which are numerically expensive due to the large system sizes required to approximate the infinite system limit. The lattice constant is set to unity and the interesting regime for the power-law exponent is $\alpha\in(2,4)$ [31]. In this range $\alpha$ is large enough for a well-defined thermodynamic limit but small enough to allow a finite FM ordering temperature $T_{c}>0$ by evading the Mermin-Wagner theorem [60]. We abbreviate $J_{\mathbf{r}_{j}}=J_{1}/|\mathbf{r}_{j}|^{\alpha}$.
@@ V.2 Dicke-Ising model: Ground-state superradiance
As a second application to a many-body quantum optical system we consider the Dicke-Ising model [61, 43, 62, 44]. This also provides an example where single-spin blocks appear in the diagrammatic expansion. Also note that the pre-computed diagrams of Sec. III are not needed in the following calculation, see below.
@@
Analogous to the other examples in this paper, we start from the $z$-polarized symmetry-unbroken phase and approach the boundary of the superradiant phase, see the arrow in Fig. 6(b). We detect the critical light-matter coupling strength $g_{c}$ from the divergence of the static FM correlator $G_{\mathbf{k}=\mathbf{0}}^{xx}(i\nu=0)$,
@@
Here and in the following we drop the zero-Matsubara frequency argument from all quantities. The static part of the retarded $xx$-interaction (44) is $J_{\mathbf{k}=\mathbf{0}}^{xx}=-2g^{2}/(\omega N)$ and $\Sigma^{xx}$ denotes the $J^{xx}$-irreducible part of $G^{xx}$. Next, we expand $\Sigma^{xx}$ in $J^{xx}$ and $V$ with the non-interacting Hamiltonian being $H_{0,h-V}=-\left[h-V\right]\sum_{i}S_{i}^{z}$. The first observation is that due to the limit $N\rightarrow\infty$ any occurrence of an interaction $J_{ii^{\prime}}^{xx}\sim 1/N$ needs to be accompanied with a free site summation. Since these diagrams are necessarily $J^{xx}$-reducible (c.f. the discussion at the end of Sec. II.2) they do not occur in $\Sigma^{xx}$ and only the $V$-interactions $\sim S_{i}^{z}S_{i^{\prime}}^{z}$ need to be considered. The second observation pertains to $z$-only blocks $G_{0,c,h-V}^{z...z}\left(\tau_{1},...,\tau_{k}\right)=b_{c}^{(k-1)}(\beta[h-V])$. In the limit $T\rightarrow 0$ where the spins are fully polarized, these generalized susceptibilities vanish except for $k=1$ for which $G_{0,c,h-V}^{z}\left(\tau_{1}\right)=S=1/2$.
These considerations equate the exact (and local) $\Sigma^{xx}$ to the infinite sum of diagrams shown in Fig. 6(c). The calculation proceeds without the kernel trick or any pre-computed diagrams of Sec. III since all diagrammatic objects are free of frequency loops and are easily evaluated at the required zero frequency. We only need $\Sigma^{xx(0)}=1/(2[h-V])$, see Tab. 2 for the TFIM. The fully static mixed-flavor blocks with two $S^{x}$ operators and $m$ appearances of the $S^{z}$ operator can be found via a $m$-fold derivative of $G_{0,c,h-V}^{xx}$ with respect to $h$ (which thus is also used as a source field),
@@ VI Conclusion
Via many examples and by applying a composite expansion strategy involving the inverse spatial dimension (or similar) as a small parameter, we showed the quantitative success of the diagrammatic approach if applied to models qualitatively described by the MF approximation. We argued that this is often the case in highly-connected spin models relevant for state-of-the-art many-body quantum optical experiments. We use a long-range Heisenberg model and the Dicke-Ising model, both on the square lattice, as a showcase. Moreover we provided various benchmark examples for nearest-neighbor models on the (hyper-)cubic lattice where our method yields accurate magnetic phase boundaries (both at $T>0$ for Ising and Heisenberg models and $T=0$ for TFIM). We emphasize that due to the analytic nature of our approach analytic continuation can be easily performed and lead to competitive results for the gap in the TFIM.
Future work could extend our approach to yet higher orders in $J$, a greater variety of spin models, including those with multiple spin-spin coupling constants, e.g. XXZ models [67], or with $\mathrm{SU}(N)$ symmetry for $N>2$ [68]. Also, the treatment of the symmetry-broken phase for the study of magnetization and spin-wave properties [47, 69] is within reach. Further, it would be interesting to consider analytic continuation beyond the computation of the gap to obtain spectral functions. These are routinely measured in inelastic neutron scattering on solid-state magnets [70] or, more recently via quench spectroscopy in Rydberg tweezer arrays [23]. For the static case, analytic insights offered by our approach are essential to shed new light [71] onto the puzzle of quantum-to-classical correspondence for static spin correlation functions in $d>1$ dimensions [72].
@@ Appendix A V-connected correlators
$\left\langle\mathcal{T}V(\tau_{1})...V(\tau_{n})A\right\rangle_{0,h,V\mkern-4.0mu-\mkern-1.5muc}\!=\!\left\langle\mathcal{T}V(\tau_{1})...V(\tau_{n})A\right\rangle_{0,h}\!-\!\!\!\!\sum_{p\in S_{n}}\!\sum_{j=0}^{n-1}\!\left\langle\mathcal{T}V(\tau_{p(1)})...V(\tau_{p(j)})A\right\rangle_{0,h,V\mkern-4.0mu-\mkern-1.5muc}\!\left\langle\mathcal{T}V(\tau_{p(j+1)})...V(\tau_{p(n)})\right\rangle_{0,h}.$ (49)
Here $S_{n}$ denotes the set of permutations of the imaginary time labels $\{1,2,...,n\}$. Standard connected spin correlators [with subscript $c$ first appearing in Eq. (13)] are defined in analogy to their $V$-connected counterparts in Eq. (49) by replacing each $V(\tau_{j})$ with a single spin operator and removing the external operators $A\rightarrow 1$.

### v3

@@ V.1 Long-range square lattice Heisenberg model
with $J_{1}<0$. This model has been recently studied with QMC simulations [42] which are numerically expensive due to the large system sizes required to approximate the infinite system limit. The lattice constant is set to unity and the interesting regime for the power-law exponent is $\alpha\in(2,4)$ [31]. In this range $\alpha$ is large enough for a well-defined thermodynamic limit but small enough for a finite FM ordering temperature $T_{c}>0$ evading the Mermin-Wagner theorem [61]. We abbreviate $J_{\mathbf{r}_{j}}=J_{1}/|\mathbf{r}_{j}|^{\alpha}$.
@@ V.2 Dicke-Ising model: Ground-state superradiance
As a second application to a many-body quantum optical system we consider the Dicke-Ising model [62, 43, 63, 44]. This also provides an example where single-spin blocks appear in the diagrammatic expansion. This means that the pre-computed diagrams of Sec. III cannot be used in the following calculation, see below.
@@
Analogous to the other examples in this work, we start from the $z$-polarized symmetry-unbroken phase and approach the boundary of the superradiant phase, see the arrow in Fig. 6(b). We detect the critical light-matter coupling strength $g_{c}$ from the divergence of the static FM correlator $G_{\mathbf{k}=\mathbf{0}}^{xx}(i\nu=0)$,
@@
Here and in the following we drop the zero-Matsubara frequency argument from all quantities. The static part of the retarded $xx$-interaction (44) is $J_{\mathbf{k}=\mathbf{0}}^{xx}=-2g^{2}/(\omega N)$ and $\Sigma^{xx}$ denotes the $J^{xx}$-irreducible part of $G^{xx}$. Next, we expand $\Sigma^{xx}$ in $J^{xx}$ and $V$ with the non-interacting Hamiltonian being $H_{0,h-V}=-\left[h-V\right]\sum_{i}S_{i}^{z}$. The first observation is that due to the limit $N\rightarrow\infty$ any occurrence of an interaction $J_{ii^{\prime}}^{xx}\sim 1/N$ needs to be accompanied with a free site summation. Since these diagrams are necessarily $J^{xx}$-reducible (c.f. the discussion at the end of Sec. II.2) they do not occur in $\Sigma^{xx}$ and only the $V$-interactions $\sim S_{i}^{z}S_{i^{\prime}}^{z}$ need to be considered. The second observation pertains to $z$-only blocks $G_{0,c,h-V}^{z...z}\left(\tau_{1},...,\tau_{m}\right)=b_{c}^{(m-1)}(\beta[h-V])$. In the limit $T\rightarrow 0$ where the spins are fully polarized, these generalized susceptibilities vanish except for $m=1$ for which $G_{0,c,h-V}^{z}\left(\tau_{1}\right)=S=1/2$.
These considerations equate the exact (and local) $\Sigma^{xx}$ to the infinite sum of diagrams shown in Fig. 6(c). The calculation proceeds without the kernel trick since all diagrammatic objects are free of frequency loops and are easily evaluated at the required zero frequency. We only need $\Sigma^{xx(0)}=1/(2[h-V])$, see Tab. 2 for the TFIM. The fully static mixed-flavor blocks with two $S^{x}$ operators and $m$ appearances of the $S^{z}$ operator can be found via a $m$-fold derivative of $G_{0,c,h-V}^{xx}$ with respect to $h$ (which thus is also used as a source field),
@@ VI Conclusion
Via many examples and by applying a composite expansion strategy involving the inverse spatial dimension (or similar) as a small parameter, we showed the quantitative success of the diagrammatic approach if applied to models qualitatively described by the MF approximation. We argued that this is often the case in highly-connected spin models relevant for state-of-the-art many-body quantum optical experiments. We use a long-range Heisenberg model and the Dicke-Ising model, both on the square lattice, as a showcase. Moreover we provided various benchmark examples for nearest-neighbor models on the (hyper-)cubic lattice where our method yields accurate magnetic phase boundaries (both at $T>0$ for Ising and Heisenberg models and $T=0$ for TFIM). We emphasize that due to the analytic nature of our approach, continuation to real frequencies is easily performed. For example we showed competitive results for the gap in the TFIM.
Future work could extend our approach to yet higher orders in $J$ (for which diagram creation can be automated) or a greater variety of spin models, including those with non-trivial unit cells or various forms of disorder. Also spin-spin couplings between sites $i,i^{\prime}$ that are characterized by two or more non-zero parameters in the $3\times 3$ coupling matrix $J_{ii^{\prime}}^{\gamma\gamma^{\prime}}$ [c.f. Eq. (1)] like e.g. XXZ models [68] could be studied with little extra effort. Another option would be to extend our approach to systems with $\mathrm{SU}(N)$ symmetry for $N>2$ [69, 70]. Also, the treatment of the symmetry-broken phase for the study of magnetization and spin-wave properties [48, 71] is within reach. Further, it would be interesting to consider analytic continuation beyond the computation of the gap to obtain spectral functions. These are routinely measured in inelastic neutron scattering on solid-state magnets [72] or, more recently via quench spectroscopy in Rydberg tweezer arrays [23]. For the static case, analytic insights offered by our approach proved essential to shed new light [73] onto the puzzle of quantum-to-classical correspondence for static spin correlation functions in $d>1$ dimensions [74].
@@ Appendix A V-connected correlators
$\left\langle\mathcal{T}V(\tau_{1})...V(\tau_{n})A\right\rangle_{0,h,V\mkern-4.0mu-\mkern-1.5muc}=\left\langle\mathcal{T}V(\tau_{1})...V(\tau_{n})A\right\rangle_{0,h}\!-\sum_{S\subsetneq\{1,...,n\}}\left\langle\mathcal{T}\prod_{j\in S}V(\tau_{j})A\right\rangle_{0,h,V\mkern-4.0mu-\mkern-1.5muc}\left\langle\mathcal{T}\prod_{k\in\{1,..,n\}\backslash S}V(\tau_{k})\right\rangle_{0,h}.$ (49)
Note that the ordering of (bosonic) operators behind imaginary time-ordering operator $\mathcal{T}$ does not matter. Standard connected spin correlators [with subscript $c$ first appearing in Eq. (13)] are defined in analogy to their $V$-connected counterparts in Eq. (49) by replacing each $V(\tau_{j})$ with a single spin operator and removing the external operators $A\rightarrow 1$.